OpenStudy (anonymous):

HELP*logs*!!!! asap please!! Solve 2 (^x + 3) = 7

OpenStudy (anonymous):

\[2^{x+3}=7\] take ln of both sides \[ln (2^{x+3})=ln(7)\] log rules \[(x+3)ln (2)=ln(7)\] \[(x+3)=ln(7)/ln(2)\] \[(x)=ln(7)/ln(2) -3\]

OpenStudy (anonymous):

i solved that...still didnt come up with the answer.. what result did you come up with??

OpenStudy (anonymous):

well , my last line was the answer

OpenStudy (anonymous):

can you give as a decimal

OpenStudy (anonymous):

x ≈ –0.19 x ≈ 0.94 x ≈ –5.19 x ≈ 2.81

OpenStudy (anonymous):

Dude / Dudette - you need to learn to use calculator for it.

OpenStudy (anonymous):

i am using calc, just dont know how to find as decimal

OpenStudy (anonymous):

i got it, -.19

OpenStudy (anonymous):

x approximately equals 0.19 seems to be the answer.

OpenStudy (anonymous):

can u help me with my other log problem please

OpenStudy (mathmate):

Just in case, is the question 2 (^x + 3) = 7 or is it (2^x)+3 = 7

OpenStudy (anonymous):

It is 2^(x+3)

OpenStudy (anonymous):

It is being written down in a very weird way by the person asking the question. But, that is the idea!

OpenStudy (mathmate):

Sorry, just to make sure, sometime the parentheses are added on by the poster because of the superscripts. So there is a chance of transcription error.

OpenStudy (anonymous):

Not in this case. We confirmed.

OpenStudy (mathmate):

Oh! Thanks and sorry for the interruption.